If a merchant has two types of tea, one work $2.70 per kilogram and the other worth $3.00 per kilogram, how many kilograms of each type should the merchant use in order to produce 30 kilograms of a blend that is worth $2.95 per kilogram?
Answer by mananth(16946)

Price ----------------quantity
Tea Type I 2.7----------------x pounds
Tea Type II 3----------------30-x pounds
Total 2.95----------------30
2.7x+3(30-x)=2.95*30
2.7x+90 -3x=88.5
2.7x-3x=88.5-90
-0.3x=-1.5
/-0.3
x=5 pounds$ 2.70
25 pounds $3.00